# From Many Models, One: Macroeconomic Forecasting with Reservoir Ensembles

### v1

@@ 3 Expert Ensembles
We now turn to an in-depth discussion of the Follow-the-Leader and Hedge online learning methods. Table 1 provides an overview of these schemes, as well as classical combination approaches based on averaging and rolling mean-squared-error weighting (see Appendix C for more details). In Section 3.3, we theoretically study the stochastic setting with and without temporal dependence, and derive regret guarantees for both FTL and decreasing Hedge.
@@ 3.1 Follow-the-Leader
with the convention that $\omega^{(k)}_{\textnormal{{ftl}},1}=1/K$. FTL is known to perform well in several settings, particularly in the case of i.i.d. losses. However, it has the downside of being sensitive in the case of adversarially generated losses due to overreacting to short-term fluctuations. de Rooij et al., (2014) derives an upper bound on the cumulative regret incurred by a forecaster using the FTL strategy due to the variability of the losses and the number of leader changes.
Let $s_{t}:=\max_{k\in[K]}\ell^{(k)}_{t}-\min_{k\in[K]}\ell^{(k)}_{t}$ and $S_{t}:=\max\{s_{1},\ldots,s_{t}\}$ denote the instantaneous maximal loss differential and maximum loss differential over $t$ periods, respectively. Define $c_{t}=1$ if there exists an expert $k^{\prime}$ with $k^{\prime}\in\mathrm{FTL}_{t-1}$ but $k^{\prime}\notin\mathrm{FTL}_{t}$ (i.e., the leader set changes at $t$), and $c_{t}=0$ otherwise. We also write that $c_{t}=\mathds{1}\{\textnormal{{ftl}}_{t-1}\not=\textnormal{{ftl}}_{t}\}$. Let $C_{t}:=\sum_{\tau=1}^{t}c_{\tau}$ be the total number of times the leader set changes up to time $t\in[T]$.
@@
Lemma 3.1 shows that FTL is effective whenever leader changes are infrequent and the loss range is small. Under i.i.d. losses whose expectations are well separated, frequent leader changes are unlikely (de Rooij et al.,, 2014). In Section 3.3, we provide a sharper result in a stochastic setting, while also allowing for dependence.
@@ 3.2 Hedge and Adaptive Hedge
Even though FTL relies on identifying a single best expert for prediction, the underlying setup can be generalized to non-sparse weighting vectors. The overarching framework is the Hedge family of algorithms (Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997, 1999; Cesa-Bianchi et al.,, 1997), also known as exponential weighted average forecasters. In Hedge, experts are weighted according to
@@
where $\{\eta_{t}\}_{t=1}^{T}$, with $\eta_{t}>0$ for all $t\in[T]$, is the sequence of the so-called learning rates (Cesa-Bianchi and Lugosi,, 2006). We use the convention $\omega^{(k)}_{\textnormal{{HDG}},1}=1/K$. Variations of Hedge propose different sequences of learning rates aimed at minimizing the cumulative regret of the forecaster. A natural way to see why (6) is effective is to note that it is the partial derivative of the scaled LogSumExp function of the scaled negative cumulative losses, that is
@@
where $\eta_{t}>0$ acts as an inverse temperature or smoothing parameter. Large $\eta_{t}$ values make the distribution over experts concentrate around the best expert(s) (approaching FTL), while small $\eta_{t}$ values yield nearly uniform weights. [^4] The LogSumExp is often referred to as the “RealSoftMax” function, as it is a differentiable approximation of the max operator. With a negative scaling of arguments and output as in the definition of $\mathcal{L}$, it hence approximates the minimum of a vector. This is also the solution (up to a constant) of the entropy-regularized optimization problem $\min_{\bm{\omega}\in\Delta^{K-1}}\left(\bm{\omega}^{\top}\bm{L}_{t-1}+\eta_{t}^{-1}H(\bm{\omega})\right)$, where $H(\bm{\omega})$ is the Shannon entropy. The dual formulation of this problem also involves the LogSumExp function, and both perspectives yield the same solution (6), see Boyd and Vandenberghe, (2004). From this viewpoint, Hedge solves a smoothed version of the Follow-the-Leader scheme, where $\eta_{t}$ controls the per-round smoothing of the loss.
@@
Setting $\eta_{t}\equiv\eta>0$ yields a constant-rate forecaster with a regret bound
@@
which leads to $\overline{R}_{T}\leq S_{T}\sqrt{T\log K/2}$. If losses take values in $[0,1]$, then $S_{T}\leq 1$ and optimal $\eta^{\star}=\sqrt{8\log K/T}$ (Littlestone and Warmuth,, 1994).
A constant rate allows for the standard closed-form online update via instantaneous losses:
@@
Doubling-trick Hedge.
Specifically, let phase $r$ cover rounds $t\in[2^{r-1},2^{r}-1]$, and set
$\eta_{r}=\sqrt{\frac{8\log K}{S_{T}^{2}2^{r-1}}}.$
At the start of phase $r$, the forecaster resets the cumulative losses to zero and applies Hedge with constant learning rate $\eta_{r}$ until the end of the phase. Each phase therefore behaves as if the horizon were $2^{r-1}$, and the total number of phases up to round $T$ is at most $\lfloor\log_{2}T\rfloor+1$. Summing the regret bounds from (7) over all completed phases yields $\overline{R}_{T}\leq S_{T}\sqrt{2T\log K}+\mathcal{O}(S_{T}\log K)$, which matches the $\mathcal{O}(\sqrt{T\log K})$ worst-case rate of Hedge with known $T$.
Decreasing Hedge.
Decreasing Hedge is obtained by setting $\eta_{t}=c_{0}\sqrt{\log K/t}$ for a constant $c_{0}>0$ (Auer and Chiang,, 2016). While this version of Hedge generally requires tuning $c_{0}$ for a specific learning run (as it is also the case for constant Hedge), Mourtada and Gaïffas, (2019) prove that the choice $c_{0}=2$ is worst-case optimal, without requiring further assumptions on the losses sequence. Decreasing Hedge, too, achieves a worst-case regret bound of $\mathcal{O}(\sqrt{T\log K})$ (Chernov and Zhdanov,, 2010). However, in stochastic settings where losses are well-behaved, more refined guarantees are possible. In fact, the generic constant $c_{0}=2$ may be conservative when the environment is not adversarial.
@@
AdaHedge dynamically tunes the learning rate in Hedge: It is a parameter-free method which adjusts $\eta_{t}$ based on direct measurements of past performance, so that no horizon or problem-dependent parameters are used (van Erven et al.,, 2011).
Let $\bm{\omega}_{t}\in\Delta^{K-1}$ be the weight vector at time $t$, let $\bm{\ell}_{t}$ be the loss vector of $K$ experts, and let the forecaster’s loss be $\overline{\ell}_{t}:=\bm{\omega}_{t}^{\top}\bm{\ell}_{t}$. The mix loss at time $t$ for a learning rate $\eta_{t}>0$ is
@@
(see, for example, Cesa-Bianchi and Lugosi,, 2006, Chapter 2). The approximation error $\delta_{t}:=\overline{\ell}_{t}-\overline{m}_{t}$ is the mixability gap, with cumulative gap $\nabla_{t}:=\sum_{s=1}^{t}\delta_{s}$. Consequently, the regret of the forecaster can be decomposed as
@@
with the convention that $\eta_{\textnormal{{ah}},t}=\infty$ if $\nabla_{t-1}=0$. The scheme is thus parameter-free and independent of quantities available only at terminal time $T$.
@@
If $s_{t}\leq S$ for all $t$, this gives $\overline{R}_{\textnormal{{ah}},T}\leq S\sqrt{T\log K}+S\left(\frac{4}{3}\ln K+2\right)$, matching the $\mathcal{O}(S\sqrt{T\ln K})$ worst-case rate. Additionally, the leading term depends on the empirical variability $\sum_{t=1}^{T}s_{t}^{2}$ rather than directly on $T$, so the bound automatically tightens in low-variance regimes.
@@ 3.3 Online Learning Guarantees under Dependence
Following Luo and Schapire, (2015) and Mourtada and Gaïffas, (2019), one can see that the complexity of the online forecasting problem for the forecaster following a certain scheme can be generally understood in terms of the gaps between the expected loss of the best expert with respect to the expected losses of other experts. Let the per-expert gap be defined as
@@
denote the so-called sub-optimality gap, which is the gap between the expected loss of the best expert with respect to the runner-up.
@@ 3.3.1 Follow-the-Leader
Our first result is a set of regret guarantees for FTL online combination under both independence and mixing assumptions. We apply Hoeffding-type and Bernstein-type bounds to showcase the different kinds of guarantees that can be obtained, which depend on two constants that quantify the dependence of losses over time, and that are necessary ingredients in our proofs. For the sake of space, we do not include here an additional bound obtained using a more sophisticated Bernstein inequality which can be found in Appendix B.3. Moreover, we emphasize that, even though the presented bounds assume a constant-in-time second moment for losses, they remain valid if one uses time-dependent variances or their uniform-in-time bound.
@@ 3.3.2 Bounds for Hedge
For $\chi^{(k)}_{t}:=(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})$ for all $\in[K], kk$$\not=k^{*}$, let $\{\chi^{(k)}_{t}\}_{t\in\mathbb{Z}}$ be the excess loss process with respect to the best expert.
@@
Let Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge. Let $k^{\star}\in[K]$ be the expert with the smallest expected loss among $K$ experts. Define $\widetilde{v}_{\max}:=\max_{k\neq k^{\star}}\textnormal{Var}(\chi^{(k)}_{t})$. Then, the expected regret of decreasing Hedge is finite,
@@
We do not include here an additional bound obtained using a more sophisticated Bernstein inequality, which can be found in Appendix B.3. Moreover, we emphasize that, even though the presented bounds assume a constant-in-time second moment for losses, they remain valid if one uses time-dependent variances or their uniform-in-time bound.
@@ 3.4 Which Ensemble Method is Best for Forecasting?
The choice of a combination method is non-trivial. Theoretical upper bounds for cumulative regret depend on the sub-optimality gap $\Delta$, which depends on the unknown population mean of losses. Moreover, some methods, such as non-adaptive Hedge, typically require tuning of their learning rate, which could be problematic in practice. However, since in our non-adversarial stochastic setting where mixing dependence conditions are mildly different from independence, we can rely on numerical evidence from the literature to inform our selection. According to the numerical results in Mourtada and Gaïffas, (2019), FTL, decreasing Hedge, and AdaHedge all demonstrate strong performance, both for large $\Delta$-gap scenarios and “hard” stochastic zero-gap ($\Delta=0$) settings, considerably outperforming the constant and doubling-trick Hedge. However, the simulations by van Erven et al., (2011) show that constant and decreasing Hedge perform similarly in practice. Balancing empirical performance with simplicity, we focus on two methods for our empirical analysis in Section 5. We implement constant Hedge as a computationally simple baseline exponential combination method, decreasing Hedge with theoretically optimal scaling ($c_{0}=2$) and AdaHedge, which does not depend on any tuning parameter.
@@
In practice, one departs from the standard online learning setting in two relevant ways. First, the so-called “forecaster’s loss”, $\overline{\ell}_{t}:=\bm{\omega}_{t}^{\top}\bm{\ell}_{t}$, is a convex combination of expert losses, whereas the performance is often measured using the loss of the combined forecast, $\ell(\widehat{Y}_{t+h},Y_{t+h})$. If $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a convex map, then $\ell(\widehat{Y}_{t+h},Y_{t+h})\leq\overline{\ell}_{t}$, hence FTL and Hedge regret bounds remain informative. Second, we have adopted the standard assumption that losses are bounded, $\sup_{t\geq 1}\max_{k\in[K]}\lvert\ell^{(k)}_{t}\rvert\leq 1$. In applications, this requirement is typically mild, since unless one works with heavy-tailed data, an appropriate normalization and a thresholding upper bound can be applied to unbounded losses, such as MSE or Huber loss.
@@ 4.1 Echo State Networks in Brief
The randomly drawn state parameters $A$, $C$, and $\zeta$} in (9) are constructed by first sampling $\widetilde{A}$, $\widetilde{C}$, and $\widetilde{{\zeta}}$ from appropriately chosen distributions, and then normalizing as follows:
@@
As is common in the ESN literature, we consider the ridge regression estimator for $\theta$. More precisely, let $X:=(X_{1},\ldots,X_{T-1})^{\top}\in\mathbb{R}^{(T-1)\times D}$ and $Y:=(Y_{2},\ldots,Y_{T})^{\top}\in\mathbb{R}^{T-1}$. Define
$M=\big(\mathbb{I}_{T-1}-\frac{1}{T-1}\mathbf{1}_{T-1}\mathbf{1}_{T-1}^{\top}\big),\quad\widetilde{Y}=MY,\quad\widetilde{X}=MX.$
Then the ridge regression estimator for $\theta$ is given by
$\widehat{W}_{\lambda}$ $=\arg\min_{W\in\mathbb{R}^{D}}\sum_{t=1}^{T-1}\left\|\widetilde{Y}_{t+1}-W^{\top}\widetilde{X}_{t}\right\|_{2}^{2}+\lambda\left\|W\right\|_{2}^{2}=\big(\widetilde{X}^{\top}\widetilde{X}+\lambda\mathbb{I}_{D}\big)^{-1}\widetilde{X}^{\top}\widetilde{Y}$,
$\widehat{b}_{\lambda}$ $=\frac{1}{T-1}{\bf{1}}_{T-1}^{\top}\left(Y-X\widehat{W}_{\lambda}^{\top}\right),$ (13)
where $\lambda\in\mathbb{R}^{+}$ is the regularization hyperparameter controlling the amount of shrinkage applied to the $W$ estimates. In practice, to select $\lambda$ we rely on time-series-adapted cross-validation (CV) (see, for example, Bergmeir and Benitez,, 2012; Bergmeir et al.,, 2018; Hyndman and Athanasopoulos,, 2018).

### v2

@@ 3 Expert Ensembles
We now turn to an in-depth discussion of the Follow-the-Leader and Hedge online learning methods. Table 1 provides an overview of these schemes, as well as classical combination approaches based on averaging and rolling mean-squared-error weighting (see Appendix D for more details). In Section 3.3, we theoretically study the stochastic setting with and without temporal dependence, and derive regret guarantees for both FTL and decreasing Hedge.
@@ 3.1 Follow-the-Leader
with the convention that $\omega^{(k)}_{\textnormal{{ftl}},1}=1/K$. FTL is known to perform well in several settings, particularly in the case of i.i.d. losses. However, it has the downside of being sensitive to adversarially-generated losses, as it overreacts to short-term fluctuations. de Rooij et al., (2014) derive an upper bound on the cumulative regret incurred by a forecaster using the FTL strategy due to the variability of the losses and the number of leader changes.
Let $s_{t}:=\max_{k\in[K]}\ell^{(k)}_{t}-\min_{k\in[K]}\ell^{(k)}_{t}$ and $S_{t}:=\max\{s_{1},\ldots,s_{t}\}$ denote the instantaneous maximal loss differential and maximum loss differential over $t$ periods, respectively. Define $c_{t}=1$ if there exists an expert $k^{\prime}$ with $k^{\prime}\in\mathrm{FTL}_{t-1}$ but $k^{\prime}\notin\mathrm{FTL}_{t}$ (i.e., the leader set changes at $t$), and $c_{t}=0$ otherwise. We can write $c_{t}=\mathds{1}\{\textnormal{{ftl}}_{t-1}\not=\textnormal{{ftl}}_{t}\}$. Let then $C_{t}:=\sum_{\tau=1}^{t}c_{\tau}$ be the total number of times the leader set changes up to time $t\in[T]$.
@@
Lemma 3.1 shows that FTL is effective whenever leader changes are infrequent, and the loss range is small. Under i.i.d. losses whose expectations are well separated, frequent leader changes are unlikely (de Rooij et al.,, 2014). In Section 3.3, we provide a sharper result in a stochastic setting, while also allowing for dependence.
@@ 3.2 Hedge and Adaptive Hedge
Even though FTL relies on identifying a single best expert for prediction, the underlying setup can be generalized to non-sparse weighting vectors. The overarching framework is that of the Hedge family of algorithms (Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997, 1999; Cesa-Bianchi et al.,, 1997), also known as exponentially weighted average forecasters. In Hedge, experts are weighted according to
@@
where $\{\eta_{t}\}_{t=1}^{T}$, with $\eta_{t}>0$ for all $t\in[T]$, is the sequence of the so-called learning rates (Cesa-Bianchi and Lugosi,, 2006). We use the standard convention $\omega^{(k)}_{\textnormal{{HDG}},1}=1/K$. Variations of Hedge propose different sequences of learning rates aimed at minimizing the cumulative regret of the forecaster. A natural way to interpret (6) is to note that it corresponds to the gradient of a scaled LogSumExp function of the negative cumulative losses
@@
where $\eta_{t}>0$ acts as a smoothing parameter. Large values of $\eta_{t}$ make the distribution over experts concentrate around the best expert(s) (recovering FTL), while small values yield nearly uniform weights. Equivalently, Hedge arises as the solution to an entropy-regularized linear optimization problem over the simplex, with Shannon entropy acting as a regularizer (see Boyd and Vandenberghe,, 2004). From this perspective, Hedge can be viewed as a smoothed version of FTL, with $\eta_{t}$ controlling the per-round smoothing of the loss.
@@
Setting $\eta_{t}:=\eta>0$ yields a constant-rate forecaster with a regret bound
@@
leading to $\overline{R}_{T}\leq S_{T}\sqrt{T\log K/2}$. If losses take values in $[0,1]$, then $S_{T}\leq 1$ and optimal $\eta^{\star}=\sqrt{8\log K/T}$ (Littlestone and Warmuth,, 1994). A constant rate allows for the standard closed-form online update via instantaneous losses:
@@
Decreasing Hedge (DecHedge).
Decreasing Hedge is obtained by setting $\eta_{t}=c_{0}\sqrt{\log K/t}$ for a constant $c_{0}>0$ (Auer and Chiang,, 2016). While this version of Hedge generally requires tuning $c_{0}$ for a specific learning run (as it is also the case for constant Hedge), Mourtada and Gaïffas, (2019) prove that the choice $c_{0}=2$ is worst-case optimal, without requiring further assumptions on the loss sequences. Decreasing Hedge, too, achieves a worst-case regret bound of $\mathcal{O}(\sqrt{T\log K})$ (Chernov and Zhdanov,, 2010). However, in stochastic settings where losses are well-behaved, more refined guarantees are possible. In fact, the generic constant $c_{0}=2$ may be conservative when the environment is not adversarial.
@@
AdaHedge dynamically tunes $\eta_{t}$ based on past performance, and neither horizon nor additional parameters are used (van Erven et al.,, 2011).
Let $\bm{\omega}_{t}\in\Delta^{K-1}$ be the weight vector at time $t$, let $\bm{\ell}_{t}$ be the loss vector of $K$ experts, and let the forecaster’s loss be $\overline{\ell}_{t}$. The mix loss at time $t$ for a learning rate $\eta_{t}>0$ is
@@
(see, e.g., Cesa-Bianchi and Lugosi,, 2006, Chapter 2). The approximation error $\delta_{t}:=\overline{\ell}_{t}-\overline{m}_{t}$ is the mixability gap, with cumulative gap $\nabla_{t}:=\sum_{s=1}^{t}\delta_{s}$. Consequently, the regret of the forecaster can be decomposed as
@@
with the convention that $\eta_{\textnormal{{ah}},t}=\infty$ if $\nabla_{t-1}=0$. The scheme is thus parameter-free and independent of quantities that are available only at terminal time $T$.
@@
If $s_{t}\leq S$ for all $t$, this gives $\overline{R}_{\textnormal{{ah}},T}\leq S\sqrt{T\log K}+S\left(\frac{4}{3}\ln K+2\right)$, matching the $\mathcal{O}(S\sqrt{T\ln K})$ worst-case rate. Additionally, the leading term depends on the empirical variability $\sum_{t=1}^{T}s_{t}^{2}$ rather than directly on $T$, so the bound tightens in low-variance regimes.
Finally, for completeness, Appendix C also discusses the doubling-trick Hedge algorithm.
@@ 3.3 Online Learning Guarantees under Dependence
Following Luo and Schapire, (2015) and Mourtada and Gaïffas, (2019), one can see that the complexity of the online forecasting problem for a forecaster following a certain scheme can be generally understood in terms of the gaps between the expected loss of the best expert and the expected losses of other experts. Let the per-expert gap be defined as
@@
denote the so-called sub-optimality gap between the expected loss of the best expert with respect to the runner-up.
@@ 3.3.1 Follow-the-Leader
Our first result is a set of regret guarantees for FTL online combination under both independence and mixing assumptions. We apply Hoeffding-type and Bernstein-type bounds to showcase the different kinds of guarantees that can be obtained, which depend on two constants that quantify the dependence of losses over time, and that are necessary ingredients in our proofs. For the sake of space, we do not include here an additional bound obtained using a Bernstein inequality which can be found in Appendix B.3. We emphasize that, even though the presented bounds assume a constant-in-time second moment for losses, they remain valid if one uses time-dependent variances or their uniform-in-time bound.
@@ 3.3.2 Bounds for Hedge
Let $\chi^{(k)}_{t}:=(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})$ for each $\in[K], kk$$\not=k^{\star}$. The process $\{\chi^{(k)}_{t}\}_{t\in\mathbb{Z}}$ is the excess loss of expert $k$ relative to the best expert $k^{\star}$.
@@
Let Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge. Let $k^{\star}\in[K]$ be the expert with the smallest expected loss. Define $\widetilde{v}_{\max}:=\max_{k\neq k^{\star}}\textnormal{Var}(\chi^{(k)}_{t})$. Then, the expected regret of decreasing Hedge is finite,
@@
An additional but more complex bound can be constructed using a different Bernstein inequality, which, for clarity of exposition, we derive in Appendix B.3. Again, our Hedge regret bounds are valid for time-dependent variance or for time-independent bounds thereof.
@@ 3.4 Which Ensemble Method is Best for Forecasting?
The choice of a combination method is non-trivial. Theoretical upper bounds for cumulative regret depend on the sub-optimality gap $\Delta$, which depends on the unknown population mean of losses. Moreover, some methods, such as non-adaptive Hedge, typically require tuning of their learning rate, which could be problematic in practice. However, since in our non-adversarial stochastic setting where mixing dependence conditions are mildly different from independence, we can rely on numerical evidence from the literature to inform our selection. According to the numerical results in Mourtada and Gaïffas, (2019), FTL, decreasing Hedge, and AdaHedge all demonstrate strong performance, both for large $\Delta$-gap scenarios and “hard” stochastic zero-gap ($\Delta=0$) settings, considerably outperforming the constant and doubling-trick Hedge. However, the simulations by van Erven et al., (2011) show that constant and decreasing Hedge perform similarly in practice. In Section 5, we implement constant Hedge as a computationally simple baseline exponential combination method, decreasing Hedge with theoretically optimal scaling ($c_{0}=2$) and AdaHedge, which does not depend on any tuning parameter.
@@
In practice, one departs from the standard online learning setting in two relevant ways. First, the forecaster’s loss $\overline{\ell}_{t}:=\bm{\omega}_{t}^{\top}\bm{\ell}_{t}$ is a convex combination of expert losses, whereas the performance is often measured using the loss of the combined forecast, $\ell(\widehat{Y}_{t+h},Y_{t+h})$. If $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a convex map, then $\ell(\widehat{Y}_{t+h},Y_{t+h})\leq\overline{\ell}_{t}$, hence FTL and Hedge regret bounds remain informative. Second, we have adopted the standard assumption that losses are bounded, $\sup_{t\geq 1}\max_{k\in[K]}\lvert\ell^{(k)}_{t}\rvert\leq 1$. In applications, this requirement is typically mild, since, unless one works with heavy-tailed data, an appropriate normalization and a thresholding upper bound can be applied to unbounded losses, such as MSE or Huber loss.
@@ 4.1 Echo State Networks in Brief
The randomly drawn parameters $A$, $C$, and $\zeta$} in (9) are constructed by first sampling $\widetilde{A}$, $\widetilde{C}$, and $\widetilde{{\zeta}}$ from appropriately chosen distributions, and then normalizing as follows:
@@
As is common in the ESN literature, we consider the ridge regression estimator for $\theta$. More precisely, let $X:=(X_{1},\ldots,X_{T-1})^{\top}\in\mathbb{R}^{(T-1)\times D}$ and $Y:=(Y_{2},\ldots,Y_{T})^{\top}\in\mathbb{R}^{T-1}$. Define $M=\big(\mathbb{I}_{T-1}-\frac{1}{T-1}\mathbf{1}_{T-1}\mathbf{1}_{T-1}^{\top}\big)$, $\widetilde{Y}=MY$, and $\widetilde{X}=MX.$ Then the ridge regression estimator for $\theta$ is given by
$\widehat{W}_{\lambda}$ $=\arg\min_{W\in\mathbb{R}^{D}}\{\left\|\widetilde{Y}-\widetilde{X}W\right\|_{2}^{2}+\lambda\left\|W\right\|_{2}^{2}\}=\big(\widetilde{X}^{\top}\widetilde{X}+\lambda\mathbb{I}_{D}\big)^{-1}\widetilde{X}^{\top}\widetilde{Y}$, (13)
$\widehat{b}_{\lambda}$ $=\frac{1}{T-1}{\bf{1}}_{T-1}^{\top}\left(Y-X\widehat{W}_{\lambda}\right),$ (14)
where $\lambda\in\mathbb{R}^{+}$ is the regularization hyperparameter. In practice, to select $\lambda$ we rely on time-series-adapted cross-validation (CV) (see, e.g., Bergmeir and Benitez,, 2012; Bergmeir et al.,, 2018; Hyndman and Athanasopoulos,, 2018).